OpenStudy (anonymous):

A football is kicked toward the goal. The height of the ball is modeled by the function h(x) = -16t^2 + 64t where t equals the time in seconds and h(x) represents the height of the ball at time t seconds. What is the axis of symmetry and what does it represent?

OpenStudy (anonymous):

A. x = 2; it takes 2 seconds to reach the maximum height and 2 seconds to fall back to the ground B. x = 2; it takes 2 seconds to reach the maximum height and 4 seconds to fall back to the ground C. x = 4; it takes 4 seconds to reach the maximum height and 4 seconds to fall back to the ground D. x = 4; it takes 4 seconds to reach the maximum height and 8 seconds to fall back to the ground

OpenStudy (tkhunny):

Have you considered factoring? It's very elucidating.

OpenStudy (mathmale):

The given formula for h(t) is a quadratic function. Because the coefficient (-16) of the t^2 term is negative, the curve (a parabola) opens down. To determine the axis of symmetry, we need to determine the vertex of this parabola. Remember how to do that? If done correctly, the result will be the numeric value of the axis of symmetry; the equation of this axis will be t = (?).

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \(-\frac{64}{2\times (-16)}=2\)

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(t\) by \(2\)

OpenStudy (tkhunny):

-16t^2 + 64t = -16(t^2 - 4t) = -16t(t-4) It appears t = 0 and t = 4 are the times it is on the ground (h = 0).

OpenStudy (lena772):

@tHe_FiZiCx99

OpenStudy (broskishelleh):

@e.mccormick can you help

OpenStudy (broskishelleh):

@jim_thompson5910 PLease Help

OpenStudy (broskishelleh):

@Whitemonsterbunny17 HELP!

whitemonsterbunny17 (whitemonsterbunny17):

Help? With what? @Broskishelleh

OpenStudy (broskishelleh):

@Whitemonsterbunny17 What is the answer?

OpenStudy (broskishelleh):

@ganeshie8 @Hero @jdoe0001 @thomaster @Luigi0210

whitemonsterbunny17 (whitemonsterbunny17):

Read the instructions that were given above by @satellite73 @mathmale and @tkhunny

OpenStudy (broskishelleh):

@ganeshie8 @Hero @jdoe0001 @thomaster @Luigi0210

OpenStudy (broskishelleh):

@Hero how can I do this like what is the answer

whitemonsterbunny17 (whitemonsterbunny17):

We don't give out direct answers. You'll have to calculate it for yourself with the given steps. c;

OpenStudy (anonymous):

so the answer would be A? @Whitemonsterbunny17

whitemonsterbunny17 (whitemonsterbunny17):

Use the formulas that were previously given in order to solve the problem. I'm not going to give you a direct answer.

OpenStudy (anonymous):

I did, i used -b/2a to find the axis of symmetry and got 2. So Im guessing C and D are off, so its either A or B right now im finding out the x values @Whitemonsterbunny17

OpenStudy (gahm8684):

I know im late but for anyone needing the answer; it is A